OpenStudy (anonymous):

The red light from a He-Ne (helium-neon) laser has a wavelength of 633 nm. What is the frequency of the light? A. 4.7 × 10^14 Hz B. 4.7 × 10^5 Hz C. 190 Hz D. 1.9 × 10^11 Hz

OpenStudy (michele_laino):

it is simple, since we have to apply this formula: \[\Large \lambda \nu = c\]

OpenStudy (anonymous):

okay!! what do we plug in? :/

OpenStudy (michele_laino):

here is the computation: \[\Large \nu = \frac{c}{\lambda } = \frac{{3 \times {{10}^{10}}}}{{6.33 \times {{10}^{ - 7}}}}\]

OpenStudy (anonymous):

okay!! and so we get 4.739336493E16 ?

OpenStudy (michele_laino):

where I have used these values: c=3*10^10 cm/sec \lambda=633*10^-9 meters = 6.33*10^-7 meters=6.33*1^-5 cm

OpenStudy (anonymous):

ahh okie!

OpenStudy (anonymous):

and so our solution is choice A?

OpenStudy (michele_laino):

please wait, I have made an error, here is the right computation: \[\Large \nu = \frac{c}{\lambda } = \frac{{3 \times {{10}^{10}}}}{{6.33 \times {{10}^{ - 5}}}}\]

OpenStudy (anonymous):

ohh okay!! so we get 4.739336493E14? so that is choice A still? :P

OpenStudy (michele_laino):

ok!

OpenStudy (michele_laino):

that's right! It is option A

OpenStudy (anonymous):

yay!! thank you!!

OpenStudy (michele_laino):

since the wave length is: 633 nm= 6.33*10^-5 cm

OpenStudy (anonymous):

:)

OpenStudy (michele_laino):

thank you!! :)